if 3.oog of Caldly reacts Completely with excess K₂ (204. What is the theoritical Yield of kcl
Answer #1

CaCl2 + K2C2O4 = CaC2O4 + 2 KCl

Moles of CaCl2 = mass / molar mass

= 3.00 / 111

= 0.02703

Moles of KCl = 2 * moles of CaCl2

= 2* 0.02703

= 0.0541

Mass of KCl = moles * molar mass

= 0.0541 * 74.55

= 4.0332 grams

If you have any questions please comment

If you satisfied with the solution please rate it thanks